OpenStudy (pokemon23):

The product of two consecutive integers is 56.Find the the integers.

OpenStudy (pokemon23):

Is the answer 8 and 6

OpenStudy (anonymous):

no 7 and 8 or -7 and -8

OpenStudy (pokemon23):

kk

OpenStudy (pokemon23):

thanks

OpenStudy (nikvist):

question is very funny, because your rank 64.

Directrix (directrix):

Let x and x + 1 be the consecutive integers. x(x+1) = 56 x^2 + x = 56 x^2 + x – 56 =0 (x+8)(x-7) = 0 x = - 8 or x = 7 Consecutive integers: -8 and -7 OR 7 and 8

OpenStudy (pokemon23):

Thanks Direct